# TinySQL: A Progressive Text-to-SQL Dataset for Mechanistic Interpretability Research

### v1

@@ Abstract
Mechanistic interpretability research faces a gap between analyzing simple circuits in toy tasks and discovering features in large models. To bridge this gap, we propose text-to-SQL generation as an ideal task to study, as it combines the formal structure of toy tasks with real-world complexity. We introduce TinySQL, a synthetic dataset progressing from basic to advanced SQL operations, and train models ranging from 33M to 1B parameters to establish a comprehensive testbed for interpretability. We apply multiple complementary interpretability techniques, including edge attribution patching and sparse autoencoders, to identify minimal circuits and components supporting SQL generation. Our analysis reveals both the potential and limitations of current interpretability methods, showing how circuits can vary even across similar queries. Lastly, we demonstrate how mechanistic interpretability can identify flawed heuristics in models and improve synthetic dataset design. Our work provides a comprehensive framework for evaluating and advancing interpretability techniques while establishing clear boundaries for their reliable application.
@@
Philip Quirke*,1                        Clement Neo*,2                        Abir Harrasse*,1                        Dhruv Nathawani3                        Amir Abdullah4,1
@@ 1 Introduction
Our analysis combines multiple complementary interpretability techniques to study SQL generation. Using Edge Attribution Patching (EAP), we identify minimal circuits supporting SQL query generation. In parallel, our SAE analysis reveals consistent patterns in how the same base model fine-tuned on different tasks yield similar interpretable heads. Lastly, our activation patching experiments reveal the challenges of identifying consistent mechanisms even across similar queries, providing insights into the limitations of current methods.
11footnotetext: Equal contributions. Correspondence to: Philip Quirke philip@withmartian.com. 1Martian; 2Apart Research; 3Gretel.ai; 4Cynch.ai
@@
    We introduce TinySQL, a structured text-to-SQL dataset bridging toy tasks and real-world applications. By controlling task complexity along multiple dimensions, as well as releasing models of different sizes trained on different subsets of TinySQL, we provide a robust testbed for mechanistic interpretability research.
@@
    Through systematic application of multiple interpretability methods (EAP, SAE, activation patching), we demonstrate both the strengths and limitations of current techniques for circuit identification. Our analysis reveals how circuits can vary even across similar queries, providing important insights into the reliability of current interpretability methods.
@@
    We show how mechanistic interpretability can be used to identify and remove unintended shortcuts in synthetic datasets. By tracing model behavior back to specific dataset characteristics, we demonstrate a practical approach for improving synthetic dataset design through interpretability findings.
@@ 3.1 Dataset Structure
Tasks are organized into three levels of increasing SQL complexity:
@@
Query Language Variants
. For each command level, we provide three variants of increasing linguistic complexity:
@@
The complete dataset consists of seven task variants (CS1/2/3, CS1/2/3-Syn, and CS1-Nat), each containing 100,000 examples generated through a systematic data creation pipeline that ensures consistency with our design principles.
3.2 Dataset Generation Methodology: CSx, CSx_Syn
Step 2: Target Query Generation. Based on command set level, we generate SQL queries with increasing complexity. CS1 creates basic SELECT-FROM statements, selecting random schema columns. CS2 adds ORDER BY clauses (90% probability), randomly choosing ascending or descending order. CS3 introduces aggregate functions (COUNT, SUM, AVG, MIN, MAX), applied to numeric columns with an 85% probability, ensuring compatibility with data types (e.g., SUM and AVG for numeric fields).
3.3 Dataset Generation Methodology: CSx_Nat
We additionally explore a synthetic dataset generation method using Gretel AI’s Data Designer [^1] https://gretel.ai/navigator/data-designer. By defining two categorical seed columns (table_name and instruction_phrase) and several generation columns (e.g., column_names, selected_columns), we prompt Gretel’s language models to produce varied table names, column sets, and natural instructions. We then assemble them into structured SQL queries, applying post-processing rules for syntactic correctness. This approach further diversifies our text-to-SQL dataset by introducing additional schema and instruction variety. More details in App. 3.3.
@@ 3.4 Models Trained
Each model is fine-tuned on three SQL complexity levels (CS1, CS2, CS3), combined with either base or synonym versions, resulting in six variations. We use shorthand like BM1_CS1 and BM1_CS1_Syn for BM1 fine-tuned on CS1 base and synonym versions, extending similarly to other variants. This produces 18 fine-tuned checkpoints (three base models × six datasets). Refer App.D for training details, and App.F.1.1 for examples of the model input format.
@@
All three base models reliably converge on all six dataset variants. For example, fine-tuned BM1 achieves over 85% exact-match accuracy on CS3_Syn, the most challenging variant, while BM2 and BM3 exceed 98% on most tasks. See App.D and Tab.4 for details. These results show that even smaller transformer architectures can learn robust text-to-SQL generation on TinySQL, providing a strong foundation for MI analysis.
@@ 4 Experiments
Our experiments included understanding how a model trained on a different set of tasks performs on one specific task.
@@ 4.1 Mechanistic Interpretability Techniques
Activation Patching. Given a clean input $x_{c}$ and corrupted input $x_{d}$, activation patching copies an activation on the corrupted input $h_{i}(x_{d})$ from node $i$ into the clean run. The causal effect of node $i$ is measured by comparing model outputs with and without patching: $f(x_{c};h_{i}\leftarrow h_{i}(x_{d}))-f(x_{c})$, where $f$ is some metric such as the logit difference between two tokens in the output distribution. This identifies which layers are crucial for the target computation (Meng et al., 2022).
Attribution patching. Attribution patching is a linear approximation to activation patching around the corrupted input. For node $i$, instead of running the model with the patched activation, we approximate the causal effect as $(h_{i}(x_{c})-h_{i}(x_{d}))\cdot\nabla_{h_{i}}f(x_{d})$, where $\nabla_{h_{i}}$ is the gradient of the model output with respect to activation $h_{i}$. This allows computing all patches with just two forward passes and one backward pass (Nanda, 2023).
@@ 4.2 Basic SQL Generation
To understand how BM1_CS1 generates field and table names in the CS1 task, we corrupt either the instruction or schema context by replacing names with alternatives (e.g. using "table_a" in the instruction but "table_b" in the CREATE TABLE statement, and similarly for field names). If the corruption does not hurt performance, it implies that the model does not significantly depend on that part of the context.
We find that the model consistently uses table names from the schema context while taking field names from the instruction, ignoring conflicting names in the other source. This strategy works for models trained on CS1 since field names appear identically in both locations. Hence, we created our synonym variants (CSx_Syn), which require models to develop more sophisticated mechanisms for mapping between equivalent terms rather than relying on exact matches. Given these results, we now will focus on models trained on the synonym variants to study more realistic text-to-SQL behaviors.
Activation Patching.
To minimise confounding variables, we take only the queries in CS1 where the table only has three fields. We analyze BM1_CS1-Syn’s field name prediction across four prompt classes in CS1: (1-3) selecting a single field from positions 1, 2, and 3, and (4) predicting the second field after correctly generating the first in a two-field query. For each prompt, we create two corrupted versions - one with an alternative table field and another with a neutral token. We ensure that the clean and corrupted field name is always a single token. We then patch in clean activations on a corrupted run, and measure patch effectiveness using:
@@
where $\ell(\cdot)$ denotes the logit difference between the correct and incorrect tokens.
Surprisingly, we observe high variability in important attention heads, even between similar prompts within categories. A set of examples in App. E. This is especially notable given that we have already tried to reduce the variance in tasks by investigating specific prompt classes. This variability motivates our use of complementary techniques like Edge Attribution Patching and sparse autoencoders for more robust analysis.
4.3 Finding Minimal Text-to-SQL Circuits via EAP and Ablation
To identify the minimal circuits responsible for SQL generation, we analyzed how the model process components like table name, field names, and ORDER BY clauses. We generated 15 batches of 100 paired clean and corrupted prompts for each SQL feature (full set of features detailed in App. F.1.1). We then performed EAP, selecting the 10 most important edges ranked by the task-specific metric $L$ (Eq. 1), similar to the one used for activation patching, which produces scores near 1 for essential edges and near 0 for less relevant ones.
For each feature, we ran EAP four times across all combinations of semantic field and table name variations. While we obtain 10 edges per batch, we retain only the edges appear across batches consistently - requiring 100% consistency for CS1 and CS2, and 80% for CS3 due to its increased complexity. The full results appear in Figures 14, 15, and 16 in the appendix. After running EAP for all features, we obtain a set of edges per feature. We then take the set of attention heads associated with these edges across all features, and retain them during the following ablation study.
Model Dataset Heads Recovered Accuracy (%)
@@ 4.3 Finding Minimal Text-to-SQL Circuits via EAP and Ablation
To validate these findings, we conducted systematic ablation experiments across multiple model components. We first ablated all the attention heads of the model, except those associated with the edges that EAP identified. We then evaluated the Recovered Accuracy by only using the set of attention heads identified as important. We tested both mean-ablation and zero-ablation techniques, though mean-ablation proved less informative for our analysis. Figures 3, 4, and 17 demonstrate how model performance changes when selectively removing components.
The ablation results are in Table 1. EAP performed better on smaller BM1 models than on larger BM2 and BM3 models. In BM1, it identified precise attention heads critical for the task, allowing fine-grained selection. However, in larger models, too few heads were found as the top edges were often between MLP components, resulting in low recovered accuracy (5–10%). To address this, we retained entire layers whenever at least one head was marked important by EAP. This adjustment significantly improved accuracy at the cost of including additional heads.
Figure 3: For model BM2_CS1_Syn with CS1 data, accuracy improves with more unablated layers, considering only the attention heads in those layers.
MLP and Layer Ablation.
We extended our analysis with systematic ablation of Multi-Layer Perceptrons (MLPs) across different layers, measuring their impact on model performance. Additionally, we performed zero and mean ablation on full layer outputs (including attention heads and MLPs).
We find that removing the first layer consistently degraded performance across all configurations, with accuracy dropping substantially as shown in Figs.19. Removing entire layers (both attention and MLPs) caused greater performance drops than ablating individual components, as seen when compared to MLP-specific ablations (Figs. 18).
4.4 SAE Feature Analysis for determining influential features
To understand how models encode schema information, we use SAEs to identify interpretable latent features.
Figure 4: For model BM1_CS3_Syn with CS3 data, accuracy increases with more unablated attention heads.
Experimental Setup.
@@ 4.4 SAE Feature Analysis for determining influential features
We first collect activation data from each BMi_CSi_Syn model when applied to its corresponding CSi and CSi_Syn datasets for $i\in{1,2}$. To understand how models trained on complex tasks handle simpler queries, we also create a “descending complexity" dataset by applying BM1_CS2_Syn and BM1_CS3_Syn on CS1 data.
@@
For training the SAEs, we use Belrose (2024)’s package and follow the top-k regime from Gao et al. (2024) with $k=16$, an expansion factor of 4, and 6 training epochs with a learning rate of $48e-$ 4. We find that the Fraction of Unexplained Variance (FVU) remains below 0.05 across all models and datasets. Training took approximately 1 hour for BM1, 2 hours for BM2, and 6 hours for BM3, all on a single A100 GPU. While our analysis focuses on the BM1 SAEs, we also release the suite of SAEs for BM2 and BM3.
Next, we identify maximally activating SAE features over key input elements and map them to attention head outputs. For instance, we examine SAE features activating over table field tokens in generated SQL, averaged over 200 examples. For each of the top five activating SAE features per layer, we decode them back into original attention-head output, average over these outputs, and then pick the top four heads with largest magnitude activations.
Model Dataset Layer 0 Layer 1
BM1_CS1_Syn CS1 1, 8, 10, 7 13, 4, 2, 1
BM1_CS1_Syn CS1_Syn 1, 10, 8, 7 13, 14, 6, 4
BM1_CS2_Syn CS2 1, 10, 8, 14 13, 8, 4, 1
BM1_CS2_Syn CS2_Syn 1, 10, 8, 2 10, 8, 1, 13
BM1_CS3_Syn CS1 1, 8, 10, 14 6, 1, 2, 5
BM1_CS3_Syn CS2 1, 8, 10, 14 6, 1, 2, 5
BM1_CS3_Syn CS3 1, 10, 8, 14 6, 2, 1, 13
BM1_CS3_Syn CS3_Syn 1, 10, 8, 7 2, 6, 1, 13
Table 2: From SAE feature analysis, important attention heads for generating table field tokens accurately. Note the similarity in Layer 0 heads across the three models.
Results.
Table 2 presents these influential attention heads for BM1’s table field processing. Across all three command datasets, BM1_CS*_Syn consistently uses similar attention heads in Layer 0 when processing table field tokens. This consistency suggests that attention mechanisms converge similarly across different trainings and datasets.
Most notably, the BM1_CS3_Syn model employs the same heads across CS1, CS2, CS3, and CS3_Syn datasets, suggesting that the mechanism for processing table field tokens remains largely unchanged despite increasing query complexity. This finding aligns with our understanding that the main differences between these datasets lie in the choice of SQL keywords (such as GROUP BY, SUM, etc.).
For additional analysis on how the model processes table names and ORDER BY or GROUP BY tokens, see App. G, particularly Tables  5, 6 and Table 7.

### v2

@@ Abstract
Mechanistic interpretability research faces a gap between analyzing simple circuits in toy tasks and discovering features in large models. To bridge this gap, we propose text-to-SQL generation as an ideal task to study, as it combines the formal structure of toy tasks with real-world complexity. We introduce TinySQL, a synthetic dataset, progressing from basic to advanced SQL operations, and train models ranging from 33M to 1B parameters to establish a comprehensive testbed for interpretability. We apply multiple complementary interpretability techniques, including Edge Attribution Patching and Sparse Autoencoders, to identify minimal circuits and components supporting SQL generation. We compare circuits for different SQL subskills, evaluating their minimality, reliability, and identifiability. Finally, we conduct a layerwise logit lens analysis to reveal how models compose SQL queries across layers: from intent recognition to schema resolution to structured generation. Our work provides a robust framework for probing and comparing interpretability methods in a structured, progressively complex setting.
@@
Abir Harrasse*1  Philip Quirke*1thanks:  Equal contributions. Correspondence to: philip@withmartian.com.  Clement Neo*2 Dhruv Nathawani3Luke Marks1Amir Abdullah4,1 1Martian  2Apart Research  3Gretel.ai  4Cynch.ai
@@ 1 Introduction
Our analysis combines multiple complementary interpretability techniques to study SQL generation. Using Edge Attribution Patching (EAP), we identify minimal circuits supporting SQL query generation. In parallel, our SAE analysis reveals consistent patterns in how the same base model fine-tuned on different tasks yield similar interpretable heads. Lastly, our logit lens analysis reveals a layered computation pattern: early layers capture query intent, middle layers resolve schema elements, and later layers synthesize structured SQL outputs.
@@
    We introduce TinySQL, a structured text-to-SQL dataset bridging toy tasks and real-world applications. By controlling SQL complexity across five subskills and releasing models trained on different subsets of TinySQL, we provide a robust testbed for mechanistic interpretability.
@@
    We apply multiple interpretability methods (EAP and SAE) to identify circuits across SQL subskills. We compare them, show their local minimality, and validate their relevance through performance and identifiability studies.
@@
    We present a layerwise logit lens analysis revealing how models compose SQL queries: early layers identify intent, middle layers resolve schema references, and later layers integrate all elements into structured outputs.
@@ 3.1 Dataset Structure
Tasks are organized into five levels of increasing SQL complexity:
@@
[…]
  • CS5 (Join Queries) introduces multi-table queries with JOIN clauses on matching column types. Example: “List workflow templates joined with pairs on user ID ordered by score and certification” → SELECT name, city FROM workflow_templates JOIN pairs ON workflow_templates.user_id = pairs.user_id ORDER BY average_score ASC, certification ASC.
Examples of each Command Level are presented in Appendix.C.4
Query Language Variants.
For each command level, we provide three variants of increasing linguistic complexity:
@@
The complete dataset consists of nine task variants (CS1/2/3, CS1/2/3/4/5-Syn, and CS1-Nat), each containing 100,000 examples generated through a systematic data creation pipeline that ensures consistency with our design principles.
We also include in Table 9 (see Appendix 7) several statistics on our text questions noted in the work of Mitsopoulou and Koutrika (2025) that analyzes popular Text-To-SQL benchmarks.
Namely, (a) Rarity: the ratio of the rare words to content words of an NL question Read and Read (2000); (b) Lexical density: the ratio of the content words to the total words of an NL question Read and Read (2000), and (c) Flesch reading ease, the ease of reading a sentence Flesch and Gould (1949).
Our datasets have higher lexical density and rarity than most Text-To-SQL datasets including Spiderb mc2 (2023) and BirdLi et al. (2023) (see Figure 8 in Mitsopoulou and Koutrika (2025)), and lower reading ease than all of the Text-To-SQL datasets studied in the same (see Figure 10 of the same work). This is perhaps due to the synthetic nature of our datasets which have a higher concentration of table and field names, and demonstrates that our datasets present a unique challenge with respect to other publicly available resources.
For details on our dataset generation methodology, including schema design, query construction, and instruction templating for CSx, CSx_Syn, and CSx_Nat, please refer to Appendices C.2 and C.3.
@@ 3.2 Models Trained
Our datasets comprise five SQL complexity levels (CS1, CS2, CS3, CS4 and CS5), combined with either base or synonym versions for the first three variants, resulting in nine variations (CS1,CS2,CS3,CS1_Syn,CS2_Syn,CS3_Syn, CS4_Syn, CS5_Syn and CS1_NAT. We use shorthand like BM1_CS1 and BM1_CS1_Syn for BM1 fine-tuned on CS1 base and synonym versions, extending similarly to other variants. We exclude CS1_NAT from all trainings. We train BM1 on all 8 remaining variants, while we train BM2 and BM3 only on the 6 CS1 through to CS3 variants. This produces 20 fine-tuned checkpoints. Refer App.D for training and architecture details, and App.G.1.1 for examples of the model input format.
@@
All three base models reliably converge on all dataset variants. For example, fine-tuned BM1 achieves over 85% exact-match accuracy on CS3_Syn, the most challenging variant, while BM2 and BM3 exceed 98% on most tasks. See App.D and Tab.7 for details. These results show that even smaller transformer architectures can learn robust text-to-SQL generation on TinySQL, providing a strong foundation for MI analysis.
@@ 4 Experiments
Our experiments aim to uncover how language models internally represent and execute structured code generation tasks, using interpretability tools such as Edge Attribution Patching, Sparse Autoencoders, and Logit Lens analysis.
@@ 4.2 Basic SQL Generation
To probe how BM1_CS1 uses context, we replace table or field names in either the instruction or schema (e.g., "table_a" vs. "table_b"). If performance remains stable, the model likely ignores the corrupted source.
We find the model relies on schema context for table names and on the instruction for field names, an effective shortcut in CS1, where names match. To prevent this, we created synonym variants (CSx_Syn) that require mapping between equivalent terms. We now focus on these variants to study more realistic behaviors.
4.3 Finding Locally Minimal Text-to-SQL Circuits via EAP and Ablation
To identify the locally-minimal circuits responsible for SQL generation, we analyzed how the model process components like table name, field names, and ORDER BY clauses. We generated 15 batches of 100 paired clean and corrupted prompts for each SQL feature (full set of features detailed in App. G.1.1). We then performed EAP, selecting the 10 most important edges ranked by the task-specific metric $L$, which produces scores near 0 for essential edges and near 1 for less relevant ones.
@@ 4.3 Finding Locally Minimal Text-to-SQL Circuits via EAP and Ablation
For each feature, we ran EAP four times across all combinations of semantic field and table name variations. While we obtain 10 edges per batch, we retain only the edges that appear across batches consistently - requiring 100% consistency for CS1 and CS2, and 80% for CS3 due to its increased complexity. The full results appear in Figures 14, 15, and 16 in the appendix. After running EAP for all features, we obtain a set of edges per feature. We then take the set of attention heads associated with these edges across all features, and retain them during the following ablation study.
Model Dataset Heads Recovered Accuracy
@@
To validate our findings, we ablated all attention heads except those linked to EAP-identified edges and measured Recovered Accuracy. We tested both mean- and zero-ablation (the latter being more informative). Figure 17 shows how performance varies under selective component removal.
Table 1 shows that EAP performed best on smaller BM1 models, identifying precise task-relevant heads. In larger models (BM2/BM3), top edges often involved MLPs, leading to few head selections and low recovered accuracy (5–10%). To compensate, we retained entire layers when any head was marked important, improving accuracy at the cost of granularity. In the next section, we address this limitation using more targeted SAE circuits.
4.4 Sparse Circuit Identification: Method and Results
To identify interpretable latent features relevant to SQL task performance, we train Sparse Autoencoders (SAEs) on attention head output and MLP activations from the decoder blocks of our transformer models. We focus our main analysis on the BM1 family spanning all the datasets, with results for BM2 provided in Appendix.I.
Experimental Setup
@@ 4.4 Sparse Circuit Identification: Method and Results
We begin by collecting attention heads and MLP activations from each of $\texttt{BM}_{2}\_\texttt{CS}_{1}\_\texttt{Syn}$ and $\texttt{BM}_{1}\_\texttt{CS}_{i}\_\texttt{Syn}$ models when applied to its corresponding $\texttt{CS}_{i}$ datasets, for $i\in\{1,3,4,5\}$ To analyze how models trained on more complex tasks generalize to simpler ones, we also create a descending complexity evaluation set by applying BM1_CS3_Syn to CS1 and CS2 data.
@@
We train SAEs on the collected activations using the Belrose (2024)’s package, following the top-$k$ training regime Gao et al. (2024) , with $k=16$, an expansion factor of 4, and a learning rate of $8\times 10^{-4}$ for 6 training epochs. Across all models and datasets, we find that the Fraction of Unexplained Variance (FVU) remains below 0.05, indicating that the SAEs faithfully reconstruct the input representations. Training each SAE takes approximately 1 hour for BM1 and 2 hours for BM2, all on a single A100 GPU.
Feature Selection via AUC Scoring.
To isolate task-relevant features, we follow the four-step procedure in Algorithm 1. We evaluate each circuit’s fidelity using zero-ablation and report recovered accuracy in Table 2, with additional results on other combinations in Appendix I.
Algorithm 1 SAE-Based Circuit Identification via Similarity-Guided Feature Selection and AUC Mapping
1:  Collect Activations: Extract residual activations $\{x^{(j)}\}_{j=1}^{N}$ at hook $h$ and position $n$
2:  Compute SAE Feature Activations: Project activations through SAE encoder to obtain sparse codes $\{z^{(j)}\}$
3:  Similarity-Guided Feature Selection: Iterate $k=1$ to $K_{\max}$: compute diff vector $\Delta_{k}$ by ablating lower-ranked features, apply $\Delta_{k}$ to the model, and measure output similarity. Stop when similarity $\geq$ threshold or improvement flattens. Set $k^{\star}=k$
4:  Map to Model Components via AUC Attribution: Let $a_{f}=\lVert W_{\text{dec}}[f]\rVert$ for each selected feature $f\in F$, and $A=\sum_{f\in F}a_{f}$. Select smallest $m$ such that:
$m^{\star}=\operatorname*{argmin}_{m}\left|\frac{\sum_{i=1}^{m}a_{i}}{A}-\tau\right|$
Model Dataset Heads and MLPs Recovered Accuracy
Layer 0 Layer 1
BM1_CS1_Syn CS1 13, 8, 1, 10, 15, 3, 14, 9, 11 80.64%
All MLP neurons
BM1_CS3_Syn CS1 13, 8, 1, 15, 10, 12, 14, 7, 9, 3 85.71%
846 MLP neurons
BM1_CS3_Syn CS3 13, 8, 1, 15, 10, 12, 14, 9, 7, 3 87.67%
All MLP neurons
Table 2: Recovered accuracy and layer-wise head selections for circuits identified via high-AUC SAE features. Each layer column is split into two rows per model (first row for attention heads and second for MLP neurons). Full results are in Appendix H.
5 Identifiability and Local Minimality of Derived Circuits
5.1 Identifiability via SAE-EAP Overlap
To assess the identifiability of circuits extracted using sparse autoencoders (SAEs), we compare them against circuits identified by Edge Attribution Patching (EAP). We quantify overlap between the two methods and contrast it with a random baseline. This comparison provides a new identifiability measure for SAE circuits.
In the following, we always use Synonyms circuits so we will omit the notation SYN.
Our results in Table 3 reveal a consistently higher overlap between EAP and SAE circuits than between EAP and random circuits (for which the accuracies are presented in Appendix.K). This supports the hypothesis that SAE-selected features align with the ground truth structure identified via causal intervention. Notably, Layer 2 exhibits more variability and lower overlap across methods, which is consistent with our earlier ablation study (Fig.19) indicating its minor role in SQL task processing, often involving only a handful of heads.
Model Dataset Layer 1 Layer 2
CS CS1 87.5% 50%
52.62%
BM1-CS3 CS1 70% 57%
74.85%
BM1-CS3 CS2 70% 50%
54.81%
BM1-CS3 CS3 50% 62.5%
36.16%
Table 3: Percentage overlap between EAP circuits and other methods across different datasets and layers. For each model and dataset, the first row shows overlap between SAE and EAP circuits, and the second row shows overlap between random (Rand) and EAP circuits
5.2 Local Minimality of Circuits
We also analyze the local minimality of the circuits—(i.e., the property that no smaller circuit can be obtained by removing or editing components of the current “locally minimal") circuits identified by SAEs, by measuring the percentage of the total model they utilize. This metric reflects how sparse and efficient the discovered subcircuits are in reproducing task behavior. Table 4 summarizes these results across different combinations of models and datasets.
Model Dataset Percentage of Model Used
BM2-CS1 CS1 69.94%
BM1-CS1 CS1 42.69%
BM1-CS3 CS1 37.95%
BM1-CS3 CS2 42.69%
BM1-CS3 CS3 43.75%
BM1-CS4 CS4 44.79%
BM1-CS5 CS5 34.83%
Table 4: Percentage of Model Components Used by SAE Circuits
We also explored an alternative circuit extraction strategy based on mean ablation . While this approach produced more minimal circuits, their effectiveness was comparable to random circuits of the same size, suggesting poor alignment with ground-truth task structure. Thus, we opted not to include mean-ablated circuits in our main results. Detailed examples and ablation configurations are provided in Appendix J.
For example, in BM1-CS3 $\rightarrow$ CS1, the found circuit recovered 88.63% accuracy while using only 22.5% of model components. Other cases show similar patterns, with best circuits retaining high performance using only  20–30% of the model.
Finally, on BM2-CS1 $\rightarrow$ CS1, focusing only on attention heads yielded circuits using just 3.17% of the model’s heads while maintaining high accuracy. However, including MLPs raised usage to 60.71%, indicating that while attention sparsity is possible, full circuit function often relies on broader model interaction.
6 Two-Phase SQL Generation: Intent First, Grounding Later
To better understand how the model constructs SQL queries, we apply LogitLens analysis Nostalgebraist (2020). We observe a clear two-phase behavior in the generation process.
In the first phase, the model focuses on identifying the intent of the query by emitting appropriate SQL keywords (e.g., SELECT, WHERE, GROUP BY). These tokens receive higher logit probabilities in the earlier layers of the model, suggesting an early commitment to the structural form of the SQL command.
Subsequently, the model transitions into a second phase where it determines the appropriate table names to include in the query. This transition is evident from the delayed rise in logit probabilities for table-related tokens, which occurs at deeper layers, indicating the model is leveraging contextual information from the prompt to ground the query in schema elements.
Figure 3: Emergence of SQL intent (via keywords) precedes table name resolution during generation in BM2CS3. LogitLens reveals elevated logit probabilities for SQL keywords early, followed by table tokens at deeper layers.
We provide additional visualizations demonstrating similar trends across other benchmarks and model configurations in Appendix L.